OpenStudy (anonymous):

Binomial Expansion. Find the term independent of x in the expansion of [(x^2) + (2/x)]^6

OpenStudy (anonymous):

What do we define as independent of x?

OpenStudy (cwrw238):

the term would not contain x

OpenStudy (anonymous):

Okay, or answer for him @cwrw238 :P

OpenStudy (cwrw238):

(r + 1)th term of ( a + b)^n = nCr a^(n-r) b^r

OpenStudy (cwrw238):

looks like the 5th term (4 + 1)th term = 6C4 (x^2)^6-4 (2/x)^4 = 15 x^4 * 16 / x^4 = 240

OpenStudy (anonymous):

how do you know it is the 5th term ?

OpenStudy (cwrw238):

good question - the power of x in x^2 and x in 2/x had to be the same . To be honest it was an inspired guess.

OpenStudy (anonymous):

i need a real way to solve this =.= thanks for the help though!

OpenStudy (cwrw238):

i'll give it some more thought .. the answer is correct though.

OpenStudy (anonymous):

yes the answer is 240!, thanks much! :)

OpenStudy (cwrw238):

OK go back to the fornula (r + 1)th term = nCr a^(n-r) b(^r for the binomial expansion of (a + b)^2 here we have x^2(n-r) and 2^2 / x^r and n = 6 so for x to be eliminated we need 2(6 - r) = r solve for r: 12 - 2r = r 12 = 3r so r = 4 therefore we need the (4 + 1)th term ie the 5th term the other way to do this would be to expand the whole expression until you reach the term where x is eliminated

OpenStudy (cwrw238):

* 4th line - it should read 2^r / x^r

OpenStudy (cwrw238):

- not that this makes any difference to the process

OpenStudy (cwrw238):

hope this is clear to you

OpenStudy (anonymous):

ok got it, thanks a bunch !

OpenStudy (cwrw238):

yw - that ones a bit tricky